OpenStudy (anonymous):

The volume of a prism is 45 yd3. What would change if the volume of the prism were measured in cubic feet? A. The volume of the prism would decrease. B. The volume of the prism would increase. C. The volume of the prism stays the same, but the number representing the volume increases. D. The volume of the prism stays the same, but the number representing the volume decreases.

jigglypuff314 (jigglypuff314):

converting units works because did never actually changes the actual quantity but makes it easier to read by giving it a different unit for example would you rather read 18000 seconds or 5 hours? anyways, there are 3 feet in a yard so 3^3 cubic ft = ___ yd^3 3^3 = 3*3*3 = ?

OpenStudy (anonymous):

so what answer is it

jigglypuff314 (jigglypuff314):

I getting to that :) first tell me what is 3*3*3 = ?

OpenStudy (anonymous):

27

jigglypuff314 (jigglypuff314):

yes :) so converting yd^3 to ft^3 you would multiply the yd^3 value by 27

OpenStudy (anonymous):

kinda confused

jigglypuff314 (jigglypuff314):

when you multiply a number by a number like 27 what would it do to that number?

OpenStudy (anonymous):

divide or subtract

jigglypuff314 (jigglypuff314):

mmmm more like if for example 27 times 3 = would the 3 increase or decrease?

OpenStudy (anonymous):

it will increase

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

sow what is the answer c or b

jigglypuff314 (jigglypuff314):

then as I said before: "converting units works because did never actually changes the actual quantity"

OpenStudy (anonymous):

so its c

jigglypuff314 (jigglypuff314):

yep :)